jesternmais:

A recipe needs 3/4 teaspoon black pepper and 1/4 red pepper. How much more black pepper is needed than red pepper for this recipe?

AZ:

Can you subtract fractions? 3/4 - 1/4

TETSXPREMEE:

whats higher

jesternmais:

yea I can I jus want to know if I got it right

AZ:

Remember when we subtract fractions, we need to have the same denominator (which we do in this question) \(\dfrac{a}{c} - \dfrac{b}{c} = \dfrac{a-b}{c}\)

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jesternmais yea I can I jus want to know if I got it right \(\color{#0cbb34}{\text{End of Quote}}\) Yeah, we can check for you. What did you get?

jesternmais:

1/2

AZ:

And that's correct!

jesternmais:

My teacher says after you do the sum with the numerator you should put it in its lowest form and that will be your answer

AZ:

And she's correct and you did put it in its lowest form!

jesternmais:

oh thanks

AZ:

You're welcome!